Henna Co. produces and sells two products, T and O. It manufactures these products in separate factories and markets them through different channels. They have no shared costs. This year, the company sold 46,000 units of each product. Sales and costs for each product follow.

Product T Product O
Sales $800,400 $800,400
Variable costs 640,320 160,080
Contribution margin 160,080 640,320
Fixed costs 32,080 512,320
Income before taxes 128,000 128,000
Income taxes (32% rate) 44,800 44,800
Net income $83,200 $83,200

2. Assume that the company expects sales of each product to decline to 29,000 units next year with no change in unit selling price. Prepare forecasted financial results for next year following the format of the contribution margin income statement as just shown with columns for each of the two products (assume a 35% tax rate). Also, assume that any loss before taxes yields a 35% tax benefit. (Round "per unit" answers to 2 decimal places. Enter losses and tax benefits, if any, as negative values.)

3. Assume that the company expects sales of each product to increase to 60,000 units next year with no change in unit selling price. Prepare forecasted financial results for next year following the format of the contribution margin income statement shown with columns for each of the two products (assume a 35% tax rate). (Round "per unit" answers to 2 decimal places.)
Answer #1

Selling price for Product T = $800,400 / 46,000 = $17.4

Selling price for Product O = $800,400 / 46,000 = $17.4

Variable costs per units for Product T = $640,320 / 46,000 = $13.92

Variable costs per unit for Product O = $160,080 / 46,000 = $3.48

2.

Product T Product O
Sales $504,600 (29,000*$17.4) $504,600 (29,000*$17.4)
Variable costs $403,680 (29,000*$13.92) $100,920 (29,000*$3.48)
Contribution margin $100,920 $403,680
Fixed costs $32,080 $512,320
Income before taxes $68,840 ($108,640)
Income taxes (35% rate) $24,094 ($68,840*35%) $38,024 ($108,640*35%)
Net income $44,746 ($70,616)

3.

Product T Product O
Sales $1,044,000 (60,000*$17.4) $1,044,000 (60,000*$17.4)
Variable costs $835,200 (60,000*$13.92) $208,800 (60,000*$3.48)
Contribution margin $208,800 $835,200
Fixed costs $32,080 $512,320
Income before taxes $176,720 $322,880
Income taxes (35% rate) $61,852 $113,008
Net income $114,868 $209,872